OpenStudy (help_people):

Create your own factorable polynomial with a GCF. Rewrite that polynomial in two other equivalent forms. Explain how each form was created.

OpenStudy (mathstudent55):

I don't understand what is the GCF of one polynomial. The GCF is the greatest common factor of 2 or more numbers or polynomials. Start with the factors: x, x + 1, x - 2 Now multiply them together. \(x(x + 1)(x - 2) \) \(= (x^2 + x)(x - 2) \) \(= x^3 - 2x^2 + x^2 - 2x\) \(=x^3 - x^2 - 2x\) Our polynomial is: \(x^3 - x^2 - 2x\) Since we created our own polynomial from factors, now we know that this polynomial factors as \(x^3 - x^2 - 2x = x(x + 1)(x - 2)\) You can combine x with x + 1, and later x with x - 2, and have two other forms of the polynomial: \(x^3 - x^2 - 2x = (x^2 + x)(x - 2)\) \(x^3 - x^2 - 2x = (x^2 - 2x)(x + 1)\)

OpenStudy (help_people):

muchos gracias thank you so so much.

OpenStudy (help_people):

^ @mathstudent55

OpenStudy (mathstudent55):

You're welcome.